OpenStudy (anonymous):

@hartnn It seems to come out as a cone, however I'm not sure if that would be the right SA formula?

hartnn (hartnn):

that seems like half of a cylinder

OpenStudy (anonymous):

Could I do SA of a full cylinder and then divide the end result by 2?

hartnn (hartnn):

sure

hartnn (hartnn):

know the formula ?

OpenStudy (anonymous):

:) A= 376.99 / 2 = 188.495

OpenStudy (anonymous):

I got C. 60pi ... IS that correct?

hartnn (hartnn):

how did u get A = 120pi ?

OpenStudy (anonymous):

No, look again dear- no 120pi.

hartnn (hartnn):

120pi = 376.00

hartnn (hartnn):

376.99

OpenStudy (anonymous):

A= 376.99 / 2 = 188.495 60pi

hartnn (hartnn):

ok, seems correct :)